OpenStudy (anonymous):

f(g(x))=(7x+4)/(-19x-4); g(x) = (2-x)/(5x+4). Find f(x).

OpenStudy (anonymous):

do u know the properties of inverse function?

OpenStudy (anonymous):

like this \[g(g^{-1}(x))=x\]

OpenStudy (anonymous):

@average_student

OpenStudy (anonymous):

yes! is that applicable in getting the f(x)?

OpenStudy (anonymous):

yes compute \[g^{-1}(x)\] then put it in the f(g(x)) equation u will have \[f(g(g^{-1}(x)))=f(x)=\frac{7g^{-1}(x)+4}{-19g^{-1}(x)-4}\]

OpenStudy (anonymous):

it's wrong when i did the checking :(

OpenStudy (anonymous):

maybe im wrong! @dpaInc

OpenStudy (anonymous):

could u plz give us the the answer to chech it tnx

OpenStudy (anonymous):

f(x) = (3x+2)/(4x-3)

OpenStudy (anonymous):

i need the solution :((

OpenStudy (anonymous):

:(((

OpenStudy (anonymous):

what do u think dpaInc ? am i right?

OpenStudy (anonymous):

@amistre64 plz check this

OpenStudy (amistre64):

hmmm, this is an undoing is it ....

OpenStudy (amistre64):

i like your inverse concept

OpenStudy (amistre64):

\[y=\frac{2-x}{5x+4}\] \[x=\frac{2-y}{5y+4}\] \[x(5y+4)=2-y\] \[5xy+4x=2-y\] \[5xy+y+4x-2=0\] \[y(5x+1)+4x-2=0\] \[y=\frac{2-4x}{5x+1}\] \[g^{-1}(x)=\frac{2-4x}{5x+1}\]

OpenStudy (amistre64):

\[f(g(g'(x)))=f(x)=\frac{7(2-4x)+4}{-19(5x+1)-4}\]

OpenStudy (amistre64):

of course we would want to make sure the domain is fair game, but thats usually a sidenote

OpenStudy (anonymous):

tnx

OpenStudy (amistre64):

and i see i mixed up the g' lol :) i put half up top and half underneath